# Evaluating Text-to-Image Synthesis with a Conditional Fréchet Distance

### v1

@@ Abstract
Evaluating text-to-image synthesis is challenging due to misalignment between established metrics and human preferences. We propose $\mathrm{cFreD}$, a metric based on the notion of Conditional Fréchet Distance that explicitly accounts for both visual fidelity and text-prompt alignment. Existing metrics such as Inception Score (IS), Fréchet Inception Distance (FID) and CLIPScore assess either image quality or image-text alignment but not both which limits their correlation with human preferences. Scoring models explicitly trained to replicate human preferences require constant updates and may not generalize to novel generation techniques or out-of-domain inputs. Through extensive experiments across multiple recently proposed text-to-image models and diverse prompt datasets, we demonstrate that $\mathrm{cFreD}$ exhibits a higher correlation with human judgments compared to statistical metrics, including metrics trained with human preferences. Our findings validate $\mathrm{cFreD}$ as a robust, future-proof metric for the systematic evaluation of text-to-image models, standardizing benchmarking in this rapidly evolving field. We release our evaluation toolkit and benchmark in the appendix.
@@ 1 Introduction
Generative models have demonstrated remarkable capabilities in text-to-image generation [53, 17, 59, 24], text-to-video generation [78, 5, 35, 37, 60], and other modalities [26, 4, 28, 47, 82, 25]. This progress has been driven by technical breakthroughs such as the development and improved understanding of Generative Adversarial Networks (GANs) [22], Variational AutoEncoders (VAEs) [32], and more recently models based on Denoising Diffusion [65] and Flow Matching [43]. Automatic evaluation metrics have played a crucial role in guiding the development and refinement of these models by offering quantitative benchmarks that compare the quality, diversity, and fidelity of various generative approaches. The most commonly used metrics include Inception Score (IS) [62], Fréchet Inception Distance (FID) [30], and CLIPScore [29].
Figure 1: Correlation of five evaluation metrics with human preferences on three benchmark datasets (Parti-Prompts, HPDv2, and COCO). Compared to FID, $\text{FD}_{\text{\tiny{DINOv2}}}$, CLIPScore, and CMMD, our proposed method (cFreD) achieves consistently higher correlation with human judgments across all datasets.
We consider that it is time to reassess evaluation metrics for text-to-image synthesis and propose $\mathrm{cFreD}$, based on a conditional formulation of Fréchet Distance. Prior work has reported that existing metrics exhibit weak correlation with human preferences [74, 76, 34, 46]. Additionally, metrics such as IS or FID, although indicative of image quality to some extent, do not measure the alignment of generated images with input text prompts, focusing solely on comparing the generated image distributions with respect to a reference set [3]. For instance, consider a dataset with two images: one of a dog and one of a cat, each paired with their corresponding prompt. A perfect text-to-image model that mistakenly swaps these mappings (i.e. generating a cat for dog prompt and vice versa) would achieve near zero FID since the overall distribution of cats and dogs is maintained, despite the misalignment with the intended prompts. We show that $\mathrm{cFreD}$ captures better image quality assessment and conditioning on input text and results in improved correlation with human preferences (see Fig. 2).
@@
  • We show that $\mathrm{cFreD}$ correlates strongly with human evaluations, outperforming previous statistical metrics on three major human preference datasets.
@@ 2 Related Work
Text-to-Image Generation Metrics. Text-to-image generation aims to synthesize images conditioned on natural language descriptions. This field has evolved significantly since the early work on GAN-based models [81, 77]. More recently, diffusion models (e.g., GLIDE [48], Imagen [61]) and autoregressive architectures (e.g., DALL-E [57]) have achieved the state-of-the-art in generating high-quality, semantically consistent images.
Evaluating the quality of generated images typically relies on metrics like the Fréchet Inception Distance (FID [30]) and Inception Score (IS) [62], which primarily assess image realism and diversity. However, these metrics often correlate poorly with human perception and do not capture alignment between generated images and their corresponding text prompt [31]. In response, several alternative metrics have been introduced. For instance, CLIPScore [29]—which leverages the joint vision-language embeddings of the CLIP model has gained popularity as a measure of semantic consistency between images and text. Other recent proposals include CMMD [31], a metric that uses CLIP embeddings combined with Maximum Mean Discrepancy (MMD) computed with a Gaussian RBF kernel, and $\text{FD}_{\text{\footnotesize{DINOv2}}}$ [67], which is based on the finding that the DINOv2-ViT-L/14 model yields more expressive feature representations for evaluation. FWD [70] is another approach that measures the similarity between generated and real images by computing the Fréchet Distance over their wavelet coefficients. Moreover, perceptual metrics such as LPIPS [83] have been used to assess image similarity, although they do not consider the semantic alignment provided by the input text.
While these methods have advanced the evaluation of text-to-image generation, they still fall short in fully capturing human preferences—particularly regarding the interplay between image quality and text–image alignment. In contrast, our proposed metric ($\mathrm{cFreD}$) integrates the text input directly into the evaluation process, thus offering a better human-aligned metric for assessing text-to-image synthesis.
Evaluating with Human Preferences. Recent research in text-to-image evaluation has focused on incorporating human preferences to assess the quality of generated images with their corresponding textual prompts [34, 76, 74, 73, 85, 84, 51, 20, 36]. In this line of work, human preference datasets are constructed where generated images are annotated with scores or rankings based on human judgments. These annotations allow the fine-tuning of vision-language models (VLMs) to predict human preferences by analyzing both the image and its accompanying text. HPS [74] directly collects data from discord servers of generated images, creating a dataset of  98k images labeled as preferred and dispreferred and uses this data to fintune a CLIP-L/14 model to maximize the similarity of preferred images with the prompt while minimizing the similarity for the dispreferred images. PickScore [34] creates a web-app and collects a a dataset of $\sim$500k images generated from various diffusion models and is used in a similar manner to train a CLIP-H/14 model to maximize the similarity of the preferred images with their prompt. ImageReward [76] collects $\sim$137k data points from DiffusionDB [71] for annotators to create preferred and dispreferred pairs. A reward model is then trained for text-to-image synthesis evaluation. HPSv2 [73] collects  108k data points from DiffusionDB [71] and COCO captions. Then, the prompts are cleaned using GPT-4 and annotators are asked to rank between images from four diffusion models. OpenCLIP-H/14 is fine-tuned to maximize the similarity of the preferred image with the prompt. Moreover, MPS [84] collects  66k preference data points from various sources [71, 54]. Prompts are refined and augmented using GPT-4. Annotators are asked to rank images across four dimensions. Finally, a multi-aspect scoring model is trained to evaluate generated images across aesthetics, alignment, and detail.
Despite the promise of these human preference–based approaches, they require extensive human-annotated datasets that are time-consuming and labor-intensive to construct. Since these methods often train on images from specific text-to-image synthesis models, they may not work well for newer models as the metric might prefer outputs for existing models. Furthermore, because these preference-trained models typically rely on real-world data, they do not effectively adapt to the model development process when researchers train on smaller datasets such as CIFAR-100 or synthetic data, nor do they easily transfer to specialized domains such as aerial or medical imaging. In contrast, $\mathrm{cFreD}$ does not require additional training or human preference data, yet it effectively incorporates textual context during evaluation and can be readily applied to new data.

### v2

# Evaluating Text-to-Image and Text-to-Video Synthesis with a Conditional Fréchet Distance

@@ Abstract
Evaluating text-to-image and text-to-video models is challenging due to a fundamental disconnect: established metrics fail to jointly measure visual quality and semantic alignment with text, leading to a poor correlation with human judgments. To address this critical issue, we propose $\mathrm{cFreD}$, a general metric based on a Conditional Fréchet Distance that unifies the assessment of visual fidelity and text-prompt consistency into a single score. Existing metrics such as Fréchet Inception Distance (FID) capture image quality but ignore text conditioning while alignment scores such as CLIPScore are insensitive to visual quality. Furthermore, learned preference models require constant retraining and are unlikely to generalize to novel architectures or out-of-distribution prompts. Through extensive experiments across multiple recently proposed text-to-image models and diverse prompt datasets, $\mathrm{cFreD}$ exhibits a higher correlation with human judgments compared to statistical metrics , including metrics trained with human preferences. Our findings validate $\mathrm{cFreD}$ as a robust, future-proof metric for the systematic evaluation of text conditioned models, standardizing benchmarking in this rapidly evolving field. We release our evaluation toolkit and benchmark here.
@@ 1 Introduction
Generative models have demonstrated remarkable capabilities in text-to-image generation [58, 18, 63, 25], text-to-video generation [83, 6, 37, 40, 64, 68], and other modalities [27, 5, 29, 52, 87, 26]. This progress has been driven by technical breakthroughs such as the development and improved understanding of Generative Adversarial Networks (GANs) [23], Variational AutoEncoders (VAEs) [34], and more recently models based on Denoising Diffusion [70] and Flow Matching [47]. Automatic evaluation metrics have played a crucial role in guiding the development and refinement of these models by offering quantitative benchmarks that compare the quality, diversity, and fidelity of various generative approaches. The most commonly used metrics include Inception Score (IS) [66], Fréchet Inception Distance (FID) [32], and CLIPScore [31].
Figure 1: Correlation of five evaluation metrics with human preferences on three benchmark datasets (GenAI-Bench, HPDv2, and COCO). Compared to FID, $\text{FD}_{\text{\tiny{DINOv2}}}$, CLIPScore, and CMMD, our proposed method (cFreD) achieves consistently higher correlation with human judgments across all datasets.
We consider that it is time to reassess evaluation metrics for text-to-image synthesis and propose $\mathrm{cFreD}$, based on a conditional formulation of Fréchet Distance. Prior work has reported that existing metrics exhibit weak correlation with human preferences [79, 81, 36, 51]. Additionally, metrics such as IS or FID, although indicative of image quality to some extent, do not measure the alignment of generated images with input text prompts, focusing solely on comparing the generated image distributions with respect to a reference set [4]. For instance, consider a dataset with two images: one of a dog and one of a cat, each paired with their corresponding prompt. A perfect text-to-image model that mistakenly swaps these mappings (i.e. generating a cat for dog prompt and vice versa) would achieve near zero FID since the overall distribution of cats and dogs is maintained, despite the misalignment with the intended prompts. We show that $\mathrm{cFreD}$ captures better image quality assessment and conditioning on input text and results in improved correlation with human preferences (see Fig. 1 and Fig. 2). $\mathrm{cFreD}$ seamlessly extends to text-to-video generation, where it matches the best metric in correlation to human preferences on recent benchmarks, confirming that the same conditional formulation scales to the temporal domain. Moreover, we demonstrate that $\mathrm{cFreD}$ behaves monotonically under severe latent-space image corruptions and prompt perturbations, whereas traditional scores such as FID can be misleading, showcasing its robustness beyond the clean-data regime.
@@
  • We show that $\mathrm{cFreD}$ correlates strongly with human evaluations, outperforming prior statistical metrics on four text-to-image human preference datasets and matching the best metric on two text-to-video human preference datasets.
@@ 2 Related Work
Text-to-Image Generation Metrics. Text-to-image synthesis has progressed from early GANs [86, 82] to today’s diffusion and autoregressive models such as GLIDE, Imagen, and DALLE [53, 65, 61]. Quality is still commonly measured with Fréchet Inception Distance (FID) and Inception Score (IS) [32, 66], which emphasize image realism/diversity but correlate only loosely with human judgments and ignore text alignment [33]. To address this, CLIP-based metrics, including CLIPScore [31], CMMD [33], and $\text{FD}_{\text{\footnotesize DINOv2}}$ [72], compare image–text pairs in a joint representation space, while alternatives such as FWD [76] (wavelet statistics) and perceptual LPIPS [88] target visual fidelity alone. Despite these refinements, existing scores still under-represent the nuanced trade-off between image quality and semantic faithfulness that guides human preference. Our proposed $\mathrm{cFreD}$ integrates the text prompt directly into the distance computation, producing evaluations that align more closely with human ratings of text-to-image outputs.
Text-to-Video Generation Metrics. Early work repurposed image-centric scores: Fréchet Video Distance (FVD) on I3D features captures visual realism but misses motion/text cues, whereas frame-wise CLIPScore [31] adds a text component yet remains motion blind. T2VBench introduces 16 controlled temporal scenarios that expose consistency failures, while EvalCrafter [50] uses 17 objective metrics to measure perceptual quality, motion, and text alignment. JEDi [51] replaces FVD’s Gaussian assumption with an MMD on JEPA embeddings, reducing sample needs , while learning-based evaluators such as VideoScore [30] align well with human judgments but demand expensive retraining, and don’t measure the alignment of the generative distribution beyond human preferences. In contrast, our $\mathrm{cFreD}$ simultaneously captures temporal and perceptual fidelity and text alignment without retraining.
Evaluating with Human Preferences. A complementary line of research replaces statistical metrics with direct human feedback, assembling large‐scale preference datasets in which annotators pick or rank images for a given prompt [36, 81, 79, 78, 90, 89, 56, 21, 39]. Early efforts such as HPS [79] and PickScore [36] crowd-sourced $\sim$100k and $\sim$500k pairwise votes, respectively, then fine-tuned CLIP (L/14 or H/14) so that “preferred” images obtain higher image–text similarity. ImageReward [81] (137k pairs), HPSv2 [78] (108k GPT-cleaned prompts), and MPS [89] (66k multi-aspect rankings) further extend scale, data sanitation, and evaluation axes; training reward models that outperform FID/IS on human correlation. While these datasets improve alignment with user preferences, they demand expensive annotation and still entangle aesthetics, realism, and semantic faithfulness. Our work pursues a lighter-weight alternative: a metric that leverages textual context without relying on human‐preference supervision. Unlike preference-trained metrics, which may overfit to current model families seen during training and struggle with small synthetic (e.g., CIFAR-100) or domain-specific data (e.g., aerial or medical imaging), $\mathrm{cFreD}$ requires no extra training and offers a reliable plug-and-play solution across diverse settings.
@@ 3 Methodology
Dataset FID $\text{{FD}}_{\text{{{DINOv2}}}}$ CLIP CMMD cFreD
$\rho^{2}$ Rank Acc. $\rho^{2}$ Rank Acc. $\rho^{2}$ Rank Acc. $\rho^{2}$ Rank Acc. $\rho^{2}$ Rank Acc.
HPDv2 [78] 0.70 86.7 0.65 86.7 0.63 15.6 0.88 80.0 0.97 91.1
Gen-AI Bench [46] 0.50 80.0 0.66 86.7 0.04 40.0 0.22 66.7 0.68 86.7
PartiPrompts [54] 0.70 - 0.70 - 0.12 - 0.54 - 0.73 -
COCO [45] 0.08 41.7 0.29 36.1 0.00 47.2 0.22 30.6 0.33 66.7
Avg. 0.50 69.5 0.58 69.8 0.20 34.3 0.47 59.1 0.68 81.5
Table 1: Correlation to human preference and rank accuracy of statistical metrics on four text-to-image datasets (HPDv2, Parti-Prompts, COCO and Gen-AI Bench). The best results are shown in bold. $\mathrm{cFreD}$ consistently achieves the highest correlation with human preference and rank accuracy across all four datasets.
3.1 Conditional Fréchet Distance
$Q_{y|x}$ $=\mathcal{N}(\mu_{y|x},\,\Sigma_{yy|x}),$ (1)
$Q_{\hat{y}|x}$ $=\mathcal{N}(\mu_{\hat{y}|x},\,\Sigma_{\hat{y}\hat{y}|x}),$
$\mathrm{cFreD}$ $=\mathbb{E}_{x}\Bigg{[}\|\mu_{y|x}-\mu_{\hat{y}|x}\|^{2}$
[…]
4 Experiment Settings
In this section, we introduce the experimental setup and report the results on two tasks. We next compare it with preference-trained metrics, variants of $\mathrm{cFreD}$ using different backbones, and the FID ++ CLIPScore combination.
Dataset FVD JEDi IS CLIP cFreD
$\rho^{2}$ Rank Acc. $\rho^{2}$ Rank Acc. $\rho^{2}$ Rank Acc. $\rho^{2}$ Rank Acc. $\rho^{2}$ Rank Acc.
T2VQA-DB [38] 0.82 86.7 0.69 82.2 0.59 73.3 0.01 57.8 0.86 91.1
EvalCrafter [50] 0.93 80.0 0.60 20.0 0.79 80.0 0.46 40.0 0.90 100
Avg. 0.88 83.35 0.65 51.1 0.69 76.7 0.24 48.9 0.88 95.6
Table 2: Correlation to human preference and rank accuracy of statistical metrics on two text-to-video datasets (T2VQA-DB and EvalCrafter). The best results are shown in bold. $\mathrm{cFreD}$ achieves the highest average correlation and rank accuracy with two datasets, demonstrating strong alignment with human judgement in text-to-video evaluation.
4.1 Experimental Setup
Dataset. To evaluate how correlated our metric is with human preferences, we need rankings of generated images from different models using the same prompt. We use three different datasets (HPDv2 [78], Gen-AI Bench [46], and PartiPrompts [54]), which contain different generated images (each from a different model) for each prompt, along with the human preferences. We took the most preferred image per prompt as the reference image, since the real image was not always rated highest by humans. To evaluate recent models, we randomly selected 1,000 prompts from the COCO train and validation sets, ensuring these prompts were not part of the HPDv2 training and testing sets. For each prompt, we generated images using nine models listed on the Arena Leaderboard [2]. The original COCO images serve as reference images in this evaluation.
Metric Comparison & Evaluation Setting. We compare $\mathrm{cFreD}$ against four statistical metrics (FID [32], $\text{FD}_{\text{\footnotesize{DINOv2}}}$ [72], CLIPScore [31], and CMMD [33]). For text-to-video, we also compare with four statistical metrics (FVD [75], JEDi [51], Inception Score, and CLIPScore). We evaluate human correlation $\rho$ for scoring perspectives and calculate the Rank Accuracy to compare ranks. $\mathrm{cFreD}$ uses DINOv2-G/14 111timm/vit_giant_patch14_dinov2.lvd142m for image embedding and the OpenCLIP ConvNext-B Text Encoder 222 convnext_base_w.laion2b_s13b_b82k_augreg for text embeddings. For text-to-video, $\mathrm{cFreD}$ uses for text embedding and FVD’s I3D for video embeddings.
4.2 Results on Text-to-Image
As shown in Table 1, $\mathrm{cFreD}$ achieves the highest correlation with human preference and the highest rank accuracy among all statistical metrics across the four benchmarks. FID and $\text{FD}_{\text{\footnotesize{DINOv2}}}$ show only moderate alignment , while the CLIP score performs poorly and CMMD provides only marginal improvements. In contrast, $\mathrm{cFreD}$ consistently outperforms all baselines, achieving the highest correlation and ranking accuracy on every dataset. We provide full per-dataset breakdowns in the supplementary material. These results suggest that $\mathrm{cFreD}$ provides a significantly more reliable and generalizable metric for human perceptual evaluation in text-to-image generation tasks.
(a) Robustness of text distortions. We progressively corrupt the caption embeddings by (i) replacing random tokens with [MASK] (MASK), (ii) randomly dropping tokens (DROP), or (iii) hallucinating new tokens with an external LLM (HALLU).
(b) Robustness of image distortions. Images are first encoded with a VQ-VAE; we then progressively add uniform noise to a growing fraction of latent codes before decoding. Here $\mathrm{cFreD}$ rises monotonically with corruption, faithfully tracking visual degradation.
Figure 3: Robustness study on the COCO. dataset We inject noise either into the text branch (a) or the image branch (b) and compare $\mathrm{cFreD}$ with the conventional FID baseline. Across both settings, $\mathrm{cFreD}$ responds consistently to quality degradation, whereas FID can be insensitive (a) or misleading (b).
4.3 Results on Text-to-Video
Tab. 2 shows the comparison of $\mathrm{cFreD}$ with four other text-to-video statistical metrics across two benchmarks: T2VQA-DB [38] and EvalCrafter [50]. $\mathrm{cFreD}$ consistently achieves the highest rank accuracy and among the highest correlations with human preference across both datasets. While FVD performs strongly in terms of correlation (avg. $\rho^{2}=0.88$), its rank accuracy is notably lower at 83.35%. Other baselines, such as JEDi, IS, and CLIP, perform worse. These results highlight the robustness and generalizability of $\mathrm{cFreD}$, demonstrating its closer alignment with human judgment in text-to-video evaluation.
4.4 Are we done with FID?
4.5 Are we done with InceptionV3?
(a) Text-to-Image Evaluation
Dataset HPSv2 MPS VQAScore cFreD
$\rho^{2}$ Rank Acc. $\rho^{2}$ Rank Acc. $\rho^{2}$ Rank Acc. $\rho^{2}$ Rank Acc.
HPDv2 [79] 0.90 88.9 0.86 86.7 0.62 80.0 0.97 91.1
Gen-AI Bench [46] 0.79 87.7 0.45 73.3 0.99 93.3 0.68 86.7
(b) Text-to-Video Evaluation
Dataset VideoScore 1.1 VisionReward cFreD
$\rho^{2}$ Rank Acc. $\rho^{2}$ Rank Acc. $\rho^{2}$ Rank Acc.
T2VQA-DB [38] 0.02 55.6 0.01 46.67 0.86 91.1
EvalCrafter [50] 0.91 20.0 0.27 60.0 0.90 100.0
Table 3: Comparison of $\mathrm{cFreD}$ (cFreD) with metrics trained using human preference across (a) text-to-image and (b) text-to-video evaluation settings. We report squared Spearman correlation ($\rho^{2}$) and rank accuracy (%) with human preference. Best results are shown in bold.
4.6 Comparison with Human Preference Trained Metrics
Tab. 3 compares $\mathrm{cFreD}$ with existing metrics trained on human preference across text-to-image and text-to-video evaluation settings. In the text-to-image domain (Table 3a), $\mathrm{cFreD}$ consistently achieves the highest correlation and rank accuracy on HPDv2 and remains competitive on Gen-AI Bench, despite not using any human preference-based training. In the text-to-video domain (Table 3b), $\mathrm{cFreD}$ outperforms all preference-trained metrics on T2VQA-DB and achieves perfect rank accuracy on EvalCrafter, despite not using any human preference training. Notably, while VideoScore [30] achieves a slightly higher correlation on EvalCrafter ($\rho^{2}=0.91$), its rank accuracy drops to 20.0%, underscoring a gap between correlation and true preference alignment. These results highlight the generality and robustness of $\mathrm{cFreD}$ across two tasks. Unlike metrics trained on preference data, $\mathrm{cFreD}$ requires no re-training and can be readily applied to specialized domains such as medical or satellite images without any human annotation or domain-specific adaptation.
5 Ablations
[…]
(a) Effect of training data.
(b) Effect of image size.
(c) Effect of model size.
(d) Effect of number of features.
(e) Effect of IN1K Zero-shot accuracy.
Effect of the size of the pre-training dataset on cFreD. Fig. 5(a) reports the partial-dependence of $\mathrm{cFreD}$ on dataset size after controlling all other covariates. Models trained on >10\,{>}10 B images set the performance ceiling: the mixed-effects fit predicts 0.040.040.100.10 higher Spearman $\rho$ than every smaller bucket (Type-II ANOVA F=206F{=}206, $p{<}10^{-170}$). All sub-billion regimes have sizable drops, with the 10M-image bucket suffering the largest penalty (–0.102 OLS; –0.077 mixed). Once these controlled effects are considered, the seemingly “strong” 10M result in the raw plot disappears and the relation between data scale and alignment becomes monotonic. These findings confirm that sheer scale, rather than subtle confounds such as architecture or resolution, remains the single most powerful driver of $\mathrm{cFreD}$ ’s agreement with human judgments, underscoring the value of massive and well-curated corpora for text-conditioned image generation metrics.
Effect of image size on cFreD. Fig. 5(b) plots the partial-dependence of $\mathrm{cFreD}$ on input crop size (shortest side, in pixels) after controlling for data scale, parameter counts, and feature width. The controlled effect is sharply negative: $\beta=-2.0\times 10^{-4}$ Spearman $\rho$ per pixel (OLS t=29t=-29, $p<10^{-100}$; mixed-effects z=4.3z=-4.3, $p<10^{-4}$), and Type-II ANOVA tags image size as a major factor (F=860F{=}860, $p<10^{-154}$). An increase in resolution from $224\!\rightarrow\!448$ is predicted to lower correlation by $\approx 0.045$; this holds even larger inputs (e.g., $\geq\!896$), incurring steeper penalties. Moderate resolutions remain competitive—Inception-scale crops land near the peak; and the DINOv2 $518$ px setting lies on the declining slope. Apparent non-monotonicities in the raw means largely vanish once covariates are held fixed, suggesting that oversized crops dilute semantically salient regions relative to their captions (higher resolution images would need more detailed captions), degrading alignment with human judgments.
Effect of ViT model size on cFreD. Fig. 5(c) plots the partial-dependence of $\mathrm{cFreD}$ on image-encoder parameter count (log scale). Controlling for data scale, crop size, and feature width reverses the naive trend: each $10{\times}$ increase in parameters decreases Spearman $\rho$ by $1.8$ points (OLS $\beta=-0.018$, t=8.0t=-8.0, $p=2{\times}10^{-15}$) and by $2.6$ points in the mixed model (z=2.5z=-2.5, $p=0.011$). Type-II ANOVA corroborates the effect (F=64F{=}64, $p<10^{-14}$). Hence, once confounds are removed, oversized ViT backbones offer no benefit, and can even hurt, alignment with human preferences, suggesting that additional capacity is unnecessary for $\mathrm{cFreD}$.
Effect of ViT Feature dimensionality on cFreD. Fig. 5(d) shows the partial-dependence of $\mathrm{cFreD}$ on ViT embedding dimensionality. After controlling for data scale, crop size, and parameter count, each additional +100+100 channels boosts Spearman $\rho$ by only +0.0025+0.0025 (OLS $\beta=2.5{\times}10^{-5}$, $t=5.1$, $p=3.6{\times}10^{-7}$; mixed $z=17.4$, $p<10^{-15}$). Gains taper beyond $\sim\!1\,\mathrm{k}$ dims but never reverse, contrary to the apparent high-dimensional downturn in the raw plot. Widths in the 1,0241,0241,2801,280 range maximize correlation, yet leaner heads (512–640 dims) lag by barely <0.01<\!0.01. Thus, modest widening is beneficial, but very large feature spaces deliver only marginal improvements relative to their computational cost.
Effect of Zero-Shot ImageNet Accuracy on cFreD.  Fig. 5(e) shows a boxplot of the correlation between $\mathrm{cFreD}$ and human preferences as a function of zero-shot ImageNet accuracy, evaluated exclusively on image-text pretrained models [59, 67, 19, 22, 80]. Higher zero-shot accuracies generally correspond to stronger correlations with human judgments, though variance exists within each accuracy level. Models with the highest accuracy (81.37%) achieve the highest median correlation, while models at the lower end (51.53%) cluster around lower correlation values. These findings suggest that zero-shot ImageNet performance can serve as a useful indicator of downstream alignment with human preferences despite considerable variability among models with similar accuracy levels.
5.1 Resilience to distortions
Here, we provide additional evidence that FID detects failures of image-text alignment under complex perturbations. For reference, it was shown in [32] that FID accurately captures image distortions under low-level image processing distortions such as Gaussian noise and Gaussian blur. Since $\mathrm{cFreD}$ is a generalization of FID, it also inherits this resilience to such distortions. Therefore, we instead focus on two harder stress tests: (i) disrupting the image in latent space and (ii) disrupting the text conditioning signal.
Image distortions.  We follow the same setting as [74] and first encode every generated image with the pretrained Cosmos tokenizer [1], randomly replace a fixed fraction of latent tokens with codes uniformly sampled from the same codebook, and then decode the corrupted latents back. This procedure preserves the global colour statistics of the training data while progressively scrambling high-level structure. As shown in Fig. 3(b), $\mathrm{cFreD}$ monotonically increases with the corruption ratio, faithfully mirroring the visual degradation. In contrast, FID decreases between 5–20% corruption, suggesting that the images have improved, before shooting up once the latent codes become overwhelmingly noisy. The non-monotonic behaviour arises because replacing a moderate number of latents can accidentally move samples closer to the marginal distribution of the real data, something FID cannot disambiguate without joint image–text statistics. $\mathrm{cFreD}$, by cross-checking against captions, correctly penalises the emerging misalignment at every corruption level.
Text distortions.  We corrupt captions fed to the text encoder while keeping images fixed, using three ways: (i) MASK: random token masking; (ii) DROP: random token deletion; and (iii) HALLU: hallucinating a portion of the caption using (LLaMA-3.1-70B-instruct [16]). The trends (Fig. 3(a)) reveal three regimes. Low corruption: $\mathrm{cFreD}$ remains well above the constant FID baseline, signalling correct alignment. Medium corruption: for MASK and DROP, the curves drift steadily downward and converge to FID, exactly as predicted by our formulation. When the text embeddings become noisy, the cross-covariance term averages out and $\mathrm{cFreD}$ collapses to the image-only score. Hallucination: hallucinated captions drive $\mathrm{cFreD}$ below FID, indicating anti-alignment between the fabricated text and the true image content. Throughout, FID fails to detect, as it only sees unchanged pixels.
6 Discussion and Conclusion
In this work, we extend and repurpose Conditional Fréchet Inception Distance ($\mathrm{cFID}$), first introduced in [71], for text-to-image and text-to-video evaluation by replacing outdated backbones (InceptionV3) with modern vision and text encoders. Our analysis shows that transformer-based encoders (e.g., ViT, DINOv2, CLIP) align significantly better with human preferences. Building on this, we introduce $\mathrm{cFreD}$, which captures semantic consistency between text and image, unlike FID or IS, and achieves higher correlation with human judgments. Our ablation further shows that performance depends on balancing model complexity, data diversity, and representation quality, not scaling the model or dataset. We advocate $\mathrm{cFreD}$ as a robust, plug-and-play metric requiring no annotations or training. By open-sourcing our implementation, we aim to promote more systematic evaluations and accelerate generative model research.
$\mathrm{cFreD}$ also offers a promising extension. $\mathrm{cFreD}$ can incorporate human preference-trained models’ encoders as backbones for hybrid evaluation, which leverages $\mathrm{cFreD}$’s robust semantic alignment and human preference data’s fine-grained insights. It can also be adapted to other conditional tasks like image captioning or audio-to-video generation. Furthermore, $\mathrm{cFreD}$ can be generalized to evaluating multi-conditional models, such as ControlNet [87], making it broadly applicable across generative tasks.
Appendix A Image and Text Backbone Models
In our experiments, we used 46 different image backbone models and 43 different text backbone models. For vision models, we employed those trained with self-supervised learning, including ViT trained on ImageNet-1k [15], ViT trained on ImageNet-21k [73], MAE [28], DINOv2 [55], MoCOv3 [7], and I-JEPA trained on both ImageNet-1k and ImageNet-22k [3]. We also incorporated image-and-text-aligned pretrained models such as CLIP [59], MetaCLIP [80], DFN-CLIP [19], OpenCLIP [9], DataComp-CLIP [22], ConvNeXT-CLIP [67], EVA02 [20], and SigLIP [85]. Additional image models included SAM-ViT [35] and Inception V3 [74]. For text models, we employed autoencoding models such as RoBERTa [49], BERT [13], ALBERT [42], ModernBERT [77], XLM-RoBERTa [12], as well as sequence-to-sequence models including FLAN-T5 [11] and T5 [60] were used. We also used the text encoders from the aforementioned image-and-text-aligned pretrained models. The complete lists of image and text backbone models used in our experiments are presented in Tables C.1 and C.2, respectively.
Appendix B Results on Text-to-Image
We provide in-depth results on three text-to-image benchmarks: HPDv2 (Tab. B.1), Parti-Prompts (Tab. B.2), and a random selection of COCO prompts (Tab. B.3).